# Port-LLM: A Port Prediction Method for Fluid Antenna based on Large Language Models

### v1

@@ III-A4 Output Projection
In the output layer, a fully connected (FC) layer is employed in conjunction with a rearrange-tensor operation to derive the final output of the model by transforming the dimensions of $\mathbf{X}_{\text{LLM}}$ in the following manner:
$\mathbf{Y}=\text{rearrange}\left(\text{FC}\left(\mathbf{X}_{\text{LLM}}\right)\right)\in\mathbb{R}^{F\times 2\times N\times M}.$ (25)
@@ III-A5 Moving Port prediction
[…]
III-B Optimization objectives
The primary objective of this study is to forecast the moving ports of the FA for future time intervals. Attaining this objective necessitates the completion of two distinct phases. Firstly, we will employ our proposed neural network model to predict the channel tables for the forthcoming $F$ time intervals, utilizing the channel tables from the preceding $T$ time intervals as input. Subsequently, we will employ Eq. (28) to calculate the moving port of the FA corresponding to the specified future $F$ time intervals. The comprehensive implementation procedure is illustrated in Fig. 3. The pseudocode for the algorithm predicting moving ports of the FA for future moments based on Port-LLM is outlined in Algorithm 1
@@ IV-A1 Dataset
To mitigate the computational demands associated with the training of our model, we adopt a SISO system for the acquisition of training datasets. In this configuration, the antenna on the BS side remains stationary, while the UE side is equipped with the FA. This FA can move in a two-dimensional plane of dimensions $10\lambda\times 20\lambda$, situated within the $y$-$z$ plane. The quantities of movable antenna ports along the $y$-axis and $z$-axis are $M=50$ and $N=100$, respectively. The densities of ports along the $y$-axis and $z$-axis are $\rho_{y}=\rho_{z}=5$. The carrier frequency $f$ utilized in this study is 39 GHz, and we employ the clustered delay line (CDL) channel model as defined by the 3rd Generation Partnership Project (3GPP) [24]. The channel model includes 37 paths, which comprise a LoS path and 36 NLoS paths. The velocity of UEs ranges from 90 km/h to 150 km/h. Each slot contains 14 OFDM symbols, and the duration of a slot is 1 ms. Each group of 50 time slots has a sampling time. The channel corresponding to this sampling moment $T_{0}$ serves as the reference channel for that group time. Furthermore, the reference channel is accessible to the UE. To enhance the quantity and diversity of the training dataset, we conducted simulations of communication channels for 10 UEs positioned in various orientations. During the simulation of each UE, we randomly selected distinct sampling time $T_{0}$ within the designated time slot. The specific values of the Root Mean Square (RMS) angular spreads of AOD, EOD, AOA, and EOA for these 10 UEs are shown in Table I. A total of 54,300 samples were collected, with 75% of the dataset allocated for the training set and the remaining 25% designated for the test set.
@@ IV-A2 Network and Training Parameters
In the simulation of our model, the forecasting period for the FA moving ports is established at $F=8$. Concurrently, the duration for the employed channel tables is also designated as $T=8$. As previously indicated, the dimension of the FA moving port table is set as $N\times M=100\times 50$. We employ the smallest version of the GPT-2 model with 768 feature dimensions and utilize only the initial $N_{L}=6$ layers of the pre-trained GPT-2 architecture. For the LoRA fine-tuning operation, the value of the reduction ratio is set as $r=4$. Furthermore, the number of heads of the multi-head attention module employed in our model is $K=8$. And the dimension is set as $d_{\text{model}}=768$ in the multi-head attention. The Adam algorithm is employed to update the parameters. The specific simulation parameters of our model can be found in Table II. In the training process of our proposed model with the warm-up aided cosine LR schedule, the learning rate of our model $\alpha_{\text{Port-LLM}}$ undergoes a linear increase from $\alpha_{\min}=4\times 10^{-6}$ to $\alpha_{\max}=1\times 10^{-3}$ in the initial 20 epochs, known as “warm-up”, followed by a cosine decrease in the final epochs, as described in Eq. (30).
$\alpha=\alpha_{\min}+\frac{1}{2}\left(\alpha_{\max}-\alpha_{\min}\right)\left(1+\cos\left(\frac{t-T_{\max}}{T-T_{\max}}\pi\right)\right),$ (30)
where $T_{\max}$ and $T$ are the number of warm-up and total epochs, respectively. This warm-up-aided cosine annealing algorithm facilitates rapid convergence of our model in the early stages and prevents it from being stuck in local optima due to high learning rates in later stages.
@@
Number of epochs $\mathcal{K}$ 500
@@ IV-A3 Baselines
To assess the efficacy of our proposed model, we conducted a comparative analysis of various model-based and deep learning-based methods for FA moving port calculations, which served as benchmarks.
@@
  • RNN [16]: RNN is a conventional neural network architecture designed for the analysis of sequential data. In our study, we substituted the pre-trained GPT-2 model integrated into our proposed model with an RNN model. The experimental setup involved the utilization of a two-layer RNN network.
  • LSTM [17]: LSTM is a specialized form of RNNs that effectively mitigate the issues of gradient vanishing and explosion that are commonly faced by conventional RNNs when handling extended sequences. In our experiment, we employed a two-layer LSTM model as a substitute for the loaded GPT-2 model integrated within our framework.
  • GRU [18]: GRU is a streamlined adaptation of the LSTM architecture. It is engineered to maintain the capacity of LSTM for managing long-term dependencies while simultaneously decreasing computational complexity and the total number of parameters within the model. Similarly, we implemented a two-layer GRU model as an alternative to the loaded GPT-2 model within our framework.
  • Transformer [26]: The Transformer is a DL model architecture that departs from conventional recurrent neural network frameworks, including RNN, LSTM, and GRU, by utilizing the attention mechanism exclusively for the processing of sequential data. Its primary advantages include the capacity for parallel training, a robust expressive capability, and an effective means of capturing long-range dependencies. In experiment, we employed a Transformer model as a substitute for the loaded GPT-2 model within our framework. Specifically, this Transformer model consists of an 8-head multi-head attention module, with an input dimension of 768 and an embedding dimension of 512.
@@ IV-A4 Performance Metrics
To assess the performance of our proposed model, the $\text{NMSE}_{\text{t}}$ is employed to quantify the disparity between the actual channel tables $\mathbf{S}$ and the predicted $\mathbf{\hat{S}}$, as described in Eq. (31). The prediction precision of our model on the test set is measured by the accuracy defined in Eq.(32).
$\text{NMSE}_{\text{t}}=10\log_{10}\left\{\mathbb{E}\left[\frac{\lVert\mathbf{\hat{S}}-\mathbf{S}\rVert^{2}}{\lVert\mathbf{S}\rVert^{2}}\right]\right\}\left(\text{dB}\right),$ (31)
@@
Furthermore, to enhance the validation of the predicted moving port of FA at the subsequent time point, we computed the NMSE between the channel associated with the predicted port and the reference channel. This metric is referred to as $\text{NMSE}_{\text{v}}$, defined as
@@
Additionally, we conducted a comparative analysis of the spectral efficiency (SE) derived from our model against the SE achieved through the Vec Prony algorithm and the MPMP algorithm, respectively. It is computed as
@@
where $N_{\text{UE}}$ is the number of UEs, $\text{SINR}_{u}$ denotes signal-to-interference-and-noise radio of the $u$-th UE.
IV-B Performance Evaluation
Fig. 4 shows the curves of $\text{NMSE}_{\text{t}}$ and $\text{NMSE}_{\text{v}}$ for the training set and test set as the number of training epochs changes during our model training process. The data illustrated in Fig. 4 reveals that the NMSE between the predicted channel tables and the actual channel tables approaches values below $-25$ dB for both the training and test datasets as the number of training epochs increases. This convergence signifies a high level of accuracy in the channel table predictions made by our proposed model. Furthermore, during the training process, the NMSE between the predicted channels associated with the moving ports of the FA and the reference channels also approaches values below $-25$ dB for both the training and test sets as the training epochs increase. This finding indicates that the mobile ports of FA, as forecasted by our model, exhibit a high degree of accuracy. By repositioning the FA to the ports predicted by our model, it is feasible to maintain the channel in a nearly unaltered state. Fig. 5 shows the accuracy of the predicted channel tables generated by our model as it varies with the number of training epochs. From the figure, it can be seen that the accuracy of the channel tables predicted by our model is approximately 94.5%.
@@ IV-B Performance Evaluation
In practical applications addressing mobility challenges, it is common for BS antennas to be configured as multi-antenna systems. In order to verify the effectiveness of our model in the MISO system, we investigated the performance of our model at $2\times 8$, $8\times 8$, and $32\times 8$ antenna configurations at the BS-side. Additionally, we compared the prediction performance and robustness of our proposed model against other neural network-based models, taking into account different configurations of BS antennas and varying UE mobility speeds.
Figure 8: The SE versus SNR, the BS has 16 antennas, the velocity of UE is 90 km/h.
Figure 9: The SE versus SNR, the BS has 16 antennas, the velocity of UE is 120 km/h.
Fig. LABEL:fig:different_models_2_8, Fig. LABEL:fig:different_models_8_8, and Fig. LABEL:fig:different_models_32_8 provide a comparative analysis of the efficacy of our model in predicting FA moving ports relative to four other neural network-based models. This evaluation was conducted across three distinct BS antenna configurations and three varying UE mobility speeds. The horizontal axis of the figures denotes 8 consecutive prediction moments, while the vertical axis represents the NMSE between the predicted channels of moving ports and the reference channels. In each scenario, UE data was collected from 10 different orientations, with each UE contributing data over 50 consecutive moments, specifically sampling at the 7-th moment within each time period. The NMSE values predicted by the model were subsequently averaged across all data. The analysis presented in these figures indicates that our model exhibits superior predictive performance, surpassing that of the Transformer-based model. In contrast, the predictive capabilities of the RNN-based, LSTM-based, and GRU-based models are relatively constrained. This deficiency can be attributed to the limited modeling capabilities of RNN-based, LSTM-based, and GRU-based architectures when addressing complex sequential challenges. Notably, the NMSE between the channel of the mobile port predicted by our model and the reference channel is approximately $-24$ dB.
Figure 10: The SE versus SNR, the BS has 16 antennas, the velocity of UE is 150 km/h.
@@
Similarly, with a BS antenna configuration of $2\times 8$ and 10 UEs considered, we conducted a comparative analysis of the SE achieved by our proposed model against the SE derived from both the Vec Prony algorithm and the MPMP algorithm. Additionally, we also evaluated the SE under the idealized scenario (“Stationary channel”) and the absence of channel prediction (“No Prediction”). As illustrated in Fig. 8, Fig. 9, and Fig. 10, optimal performance is observed under “Stationary channel” condition. Conversely, performance is significantly diminished in the “No Prediction” condition. Furthermore, the consideration of a two-dimensional (2D) region for FA movement in our model, as opposed to the one-dimensional (1D) framework utilized by the MPMP algorithm, further enhances the performance of our model. As can be observed from these figures, the SE derived from the ports predicted by our model surpasses that obtained from both the Vec Prony algorithm and the MPMP algorithm at medium and high speeds.

### v2

@@ III-A4 Output Projection
In the Output Projection module, a two-layer linear neural network is employed in conjunction with the rearrange operation to derive the final output of the model in the following manner:
$\mathbf{Y}=\text{rearrange}\left(\text{Linear}\left(\text{GELU}\left(\text{Linear}\left(\text{rearrange}\left(\mathbf{X}_{\text{LLM}}\right)\right)\right)\right)\right),$ (27)
where $\mathbf{Y}\in\mathbb{R}^{F\times 2\times N\times M}$.
@@
Figure 4: The architecture of our proposed Prompt-Port-LLM model.
@@ III-A5 Moving Port prediction
III-B The Port-LLM model with prompt fine-tuning
The Port-LLM model that utilizes prompt fine-tuning, referred to as Prompt-Port-LLM, shares a similar network architecture with the original Port-LLM model. The primary distinction between the two models lies in their fine-tuning strategies. Specifically, the original Port-LLM employs the LoRA fine-tuning approach, while the Prompt-Port-LLM incorporates the prompt fine-tuning mechanism. Furthermore, we design a specialized Prompt Encoder module to make prompts trainable.
As illustrated in Fig. 4, the dynamic prompts information associated with the Prompt-Port-LLM model encompasses the description of the model’s task, the characteristics of the dataset, and the dynamic statistical properties of the input data throughout the training process. These dynamic statistical properties are computed in real time and include the maximum and minimum values, mean, standard deviation, and median of the input data.
Dynamic prompts need to undergo processing by the tokenizer and embedder components of the loaded pre-trained LLM. The pre-trained LLM tokenizer transforms the prompt text into a discrete sequence of symbolic tokens, while the pre-trained LLM embedder subsequently converts this sequence into a dense vector representation suitable for neural network processing. During the prompt encoding phase, we design a specialized Prompt Encoder module to deeply encode the prompt vectors. This module employs a two-layer fully connected (linear) network architecture, with the hidden layer dimension set to $d_{\text{model}}$, which facilitates dynamic optimization of the prompt through the utilization of trainable parameters.
III-C Optimization objectives
@@
Forecast the channel tables for the future $F$ time intervals $\mathbf{\hat{S}}$: $\mathbf{\hat{S}}=\text{Net}_{\text{Port}-\text{LLM}}\left(\mathbf{S}\right).$ Update our proposed model by using adaptive moment estimation: $\mathcal{L}_{\text{Port}-\text{LLM}}=\frac{\lVert\mathbf{S}-\mathbf{\hat{S}}\rVert^{2}}{\lVert\mathbf{S}\rVert^{2}}.$ Obtain the predicted moving ports of FA $\mathbf{P}=\left[\mathbf{p}_{1},\cdots,\mathbf{p}_{F}\right]$ for the subsequent $F$ moments: $\begin{aligned} \mathbf{p}_{i}=\text{unravel\_index}\bigg{(}&\text{argmin}\left(\left|\mathbf{\hat{S}}_{i}-\mathbf{H}_{\text{ref}_{i}}\right|\right),\\ &\!\left(N,M\right)\bigg{)}\end{aligned}.$ Obtain the corresponding port channel $\mathbf{h}=\left[h_{1},h_{2},\cdots,h_{F}\right]^{T}$ in the actual channel tables based on the predicted ports. Validate the accuracy of the predicted ports by our model, and calculate the NMSE between the channel of the port predicted by our model and the reference channel: $\mathcal{L}_{\text{validate}}=\frac{\lVert\mathbf{h}_{\text{ref}}-\mathbf{h}\rVert^{2}}{\lVert\mathbf{h}\rVert^{2}}.$ End for
The primary objective of this study is to forecast the moving ports of the FA for future time intervals. Attaining this objective necessitates the completion of two distinct phases. Firstly, we will employ our proposed neural network model to predict the channel tables for the forthcoming $F$ time intervals, utilizing the channel tables from the preceding $T$ time intervals as input. Subsequently, we will employ Eq. (30) to calculate the moving port of the FA corresponding to the specified future $F$ time intervals. The comprehensive implementation procedure is illustrated in Fig. 5. The pseudocode for the algorithm predicting moving ports of the FA for future moments based on Port-LLM is outlined in Algorithm 1. Note that the algorithmic flow of our proposed Prompt-Port-LLM model for predicting the moving port of FA is consistent with that of the Port-LLM model.
@@ IV Numerical Results
Note that during our proposed model training and performance testing, to reduce the number of model parameters, we employ single-antenna data in the single-input single-output (SISO) setting for model training. Nevertheless, given that practical BSs are typically equipped with multiple antennas, our trained model is directly utilized for testing in the multiple-input single-output (MISO) scenario without undergoing any retraining.
IV-A Simulation Settings
@@ IV-A1 Dataset
To mitigate the computational demands associated with the training of our model, we adopt a SISO system for the acquisition of training datasets. In this configuration, the antenna on the BS side remains stationary, while the UE side is equipped with the FA. This FA can move in a two-dimensional plane of dimensions $10\lambda\times 20\lambda$, situated within the $y$-$z$ plane. The quantities of movable antenna ports along the $y$-axis and $z$-axis are $M=50$ and $N=100$, respectively. The densities of ports along the $y$-axis and $z$-axis are $\rho_{y}=\rho_{z}=5$. The carrier frequency $f$ utilized in this study is 39 GHz, and we employ the clustered delay line (CDL) channel model as defined by the 3rd generation partnership project (3GPP) [26]. The channel model includes 37 paths, which comprise a LoS path and 36 NLoS paths. The velocity of UEs ranges from 90 km/h to 150 km/h. Each slot contains 14 OFDM symbols, and the duration of a slot is 1 ms. Each group of 50 time slots has a sampling time. The channel corresponding to this sampling moment $T_{0}$ serves as the reference channel for that group time. Furthermore, the reference channel is accessible to the UE. To enhance the quantity and diversity of the training dataset, we conduct simulations of communication channels for 10 UEs positioned in various orientations. During the simulation of each UE, we randomly select a distinct sampling time $T_{0}$ within the designated time interval. The specific values of the Root Mean Square (RMS) angular spreads of AOD, EOD, AOA, and EOA for these 10 UEs are shown in Table I. A total of 54,300 samples are collected, with 75% of the dataset allocated for the training set and the remaining 25% designated for the test set.
@@ IV-A2 Network and Training Parameters
In the simulation of our model, the forecasting period for the FA moving ports is established at $F=8$. Concurrently, the duration for the employed channel tables is also designated as $T=8$. As previously indicated, the dimension of the FA moving port table is set as $N\times M=100\times 50$. We employ the smallest version of the GPT-2 model with 768 feature dimensions and utilize only the initial $N_{L}=6$ layers of the pre-trained GPT-2 architecture. Furthermore, the number of heads of the multi-head attention module employed in our model is $K=8$. And the dimension is set as $d_{\text{model}}=768$ in the multi-head attention. In the input projection module, the hidden layer dimension $d_{l}$ of the two-layer Linear neural network is set as 2048. Furthermore, in the proposed Prompt-Port-LLM model, the hidden layer dimension of the prompt encoder module is specified as 768. The Adam algorithm is employed to update the parameters. The specific simulation parameters of our model can be found in Table II. In the training process of our proposed model with the warm-up aided cosine learning rate (LR) schedule, the learning rate of our model $\alpha_{\text{Port-LLM}}$ undergoes a linear increase from $\alpha_{\min}=4\times 10^{-6}$ to $\alpha_{\max}=1\times 10^{-3}$ in the initial 100 epochs, known as “warm-up”, as described in Eq. (32).
$\alpha_{\text{Port-LLM}}=\alpha_{\min}+\left(\alpha_{\max}-\alpha_{\min}\right)\cdot\frac{t}{T_{\max}},$ (32)
where $T_{\max}=100$ and $t$ are the number of total warm-up epochs and current number of training epoch, respectively.
In the “cosine decay” phase, the LR of the model is cosine decaying. In general, the LR of our model varies as follows during our model training:
$\alpha_{\text{Port-LLM}}=\alpha_{\min}+\frac{1}{2}\left(\alpha_{\max}-\alpha_{\min}\right)\left(1+\cos\left(\frac{t-T_{\max}}{T-T_{\max}}\pi\right)\right),$ (33)
This warm-up-aided cosine annealing algorithm facilitates rapid convergence of our model in the early stages and prevents it from being stuck in local optima due to high learning rates in later stages.
@@
Number of epochs $\mathcal{K}$ 600
@@ IV-A3 Baselines
To assess the efficacy of our proposed model, we conduct a comparative analysis of various model-based and deep learning-based methods for FA moving port calculations, which served as benchmarks.
@@
  • RNN [18]: RNN is a conventional neural network architecture designed for the analysis of sequential data. In our study, we substitute the pre-trained GPT-2 model integrated into our proposed model with an RNN model. The experimental setup involve the utilization of a two-layer RNN network.
  • LSTM [19]: LSTM is a specialized form of RNNs that effectively mitigate the issues of gradient vanishing and explosion that are commonly faced by conventional RNNs when handling extended sequences. In our experiment, we employ a two-layer LSTM model as a substitute for the loaded GPT-2 model integrated within our framework.
  • GRU [20]: GRU is a streamlined adaptation of the LSTM architecture. It is engineered to maintain the capacity of LSTM for managing long-term dependencies while simultaneously decreasing computational complexity and the total number of parameters within the model. Similarly, we implement a two-layer GRU model as an alternative to the loaded GPT-2 model within our framework.
  • Transformer [28]: The Transformer is a deep learning model architecture that departs from conventional recurrent neural network frameworks, including RNN, LSTM, and GRU, by utilizing the attention mechanism exclusively for the processing of sequential data. In experiment, we employ a Transformer model as a substitute for the loaded GPT-2 model within our framework. Specifically, this Transformer model consists of an 8-head multi-head attention module, with an input dimension of 768 and an embedding dimension of 512. In addition, the hidden layer dimension of the multilayer perceptron (MLP) inside the Transformer model is set to 3072.
@@ IV-A4 Performance Metrics
Since our proposed Port-LLM-based model predicts the moving ports of the FA in two steps, the first step is to predict the channel tables, and the second step is to obtain the future moving ports of the FA based on the predicted channel tables. Therefore, during the model training process, we set several metrics to evaluate the performance of our proposed models. Specifically, Accuracy is employed to measure the precision of the model in predicting the channel table, as described in Eq. (34).
@@
$\text{NMSE}_{\text{v}}$ is used to evaluate the NMSE between the channel corresponding to the predicted moving port based on the predicted channel table of the proposed model and the reference channel. This metric is defined in Eq. (35).
@@
Additionally, we conduct a comparative analysis of the spectral efficiency (SE) derived from our model against the SE achieved through the Vec Prony algorithm and the MPMP algorithm, respectively. It is computed as
$\text{SE}=\sum_{u=1}^{N_{\text{UE}}}{\mathbb{E}}\left\{\log_{2}\left(1+\text{SINR}_{u}\right)\right\}\ \left(\text{bps/Hz}\right),$ (36)
where $N_{\text{UE}}$ is the number of UEs, $\text{SINR}_{u}$ denotes signal-to-interference-and-noise radio of the $u$-th UE. In the process of simulating and evaluating the SE of the system, we construct a DL multiuser MISO communication scenario containing 10 UEs. The simulation parameters are set as follows: signal-to-noise ratio (SNR) is tested from 0 dB to 30 dB, and the DL precoder is eigen zero-forcing (EZF) [29].
IV-B Performance Evaluation
@@
It should be further noted that the RNN-based, LSTM-based, GRU-based, and Transformer-based architectures have experienced gradient explosion during the training process. To mitigate this problem, we incorporate a LayerNorm layer and a dropout operation within the output projection module of these models. In contrast, our proposed Port-LLM and Prompt-Port-LLM models, which are based on LLMs, do not suffer from such problems. This observation further highlights the superiority of our proposed architecture based on the pre-trained GPT-2 model.
Fig. 7 illustrates the curves of the $\text{NMSE}_{\text{v}}$ between the reference channel and the channel corresponding to the moving port predicted by our proposed Port-LLM model and Prompt-Port-LLM model with the number of training epochs on both the training set and the test set during the training process. As can be seen in Fig. 7, the NMSE between the reference channel and the channel corresponding to the moving port predicted by both models ultimately converges to slightly below -30 dB. This result further shows that our proposed Port-LLM model and Prompt-Port-LLM model exhibit high accuracy in predicting the moving ports of FA at future moments.
@@
In practical applications addressing mobility challenges, it is common for BS antennas to be configured as multi-antenna systems. In order to verify the effectiveness of our models in the MISO system, we investigate the performance of our models at $2\times 8$, $8\times 8$, and $32\times 8$ antenna configurations at the BS-side. Additionally, we compare the prediction performance and robustness of our proposed models against other neural network-based models, taking into account different configurations of BS antennas and varying UE mobility speeds. It is noteworthy that when evaluating the performance of our proposed models in MISO scenarios, we directly employ the previously trained model under SISO conditions without retraining for different MISO configurations.
Figure 10: The SE versus SNR, the BS has $2\times 8$ antennas, the velocity of UE is 90 km/h.
Figure 11: The SE versus SNR, the BS has $2\times 8$ antennas, the velocity of UE is 120 km/h.
[…]
Figure 12: The SE versus SNR, the BS has $2\times 8$ antennas, the velocity of UE is 150 km/h.
@@
Similarly, with a BS antenna configuration of $2\times 8$ and 10 UEs considered, we conduct a comparative analysis of the SE achieved by our proposed Port-LLM model and Prompt-Port-LLM model against the SE derived from both the Vec Prony algorithm and the MPMP algorithm. Additionally, we also evaluat the SE under the idealized scenario (“Stationary channel”) and the absence of channel prediction (“No Prediction”). As illustrated in Fig. 10, Fig. 11, and Fig. 12, optimal performance is observed under “Stationary channel” condition. Conversely, performance is significantly diminished in the “No Prediction” condition. Furthermore, the Port-LLM model and the Prompt-Port-LLM model we proposed achieve better SE than those obtained using the MPMP algorithm and the Vec Prony algorithm. It indicates that, compared to the MPMP algorithm and the Vec Prony algorithm, both of our proposed models are more effective in accurately capturing UE mobility in the context of UE movement, resulting in more accurate predictions of the moving ports of the FA at future moments. Additionally, the consideration of a two-dimensional (2D) region for FA movement in our proposed models, as opposed to the one-dimensional (1D) framework utilized by the MPMP algorithm, further enhances the performance of our models.
TABLE III: The Accuracy, $\text{NMSE}_{\text{v}}$, network parameters, and interference time of different models.
Model Accuracy (%) $\text{NMSE}_{\text{v}}$ (dB) Network parameters ($1\times 10^{6}$) Interference time (ms)
LSTM-based 91.16 -19.64 242.46/242.46 -
RNN-based 93.15 -20.96 235.37/235.37 -
GRU-based 92.94 -20.60 240.10/240.10 -
Transformer-based 93.40 -21.01 238.12/238.12 -
Port-LLM 98.18 -23.89 232.70/313.80 4.72
Prompt-Port-LLM 99.45 -23.91 233.00/314.92 7.52
In order to assess the deployment feasibility of our proposed models in real applications, we investigate the number of training parameters for our proposed Port-LLM model, Prompt-Port-LLM model, and other comparative neural network-based models. In addition, we also investigate the inference time required for our proposed models to perform one prediction. All of our neural network-based comparison experiments are conducted on a server equipped with 1 NVIDIA GeForce RTX 3090 GPU, 1 Intel Core i9-10920X CPU and 94 GB of RAM. As illustrated in Table III, the predictive performance of our models surpasses that of the other neural network-based models. Although the model parameters of the proposed Port-LLM and Prompt-Port-LLM models exceed those of other neural network-based models, it is important to note that a portion of these parameters are frozen, resulting in the number of parameters requiring retraining that is comparable to that of other neural network models. Furthermore, both the Port-LLM and Prompt-Port-LLM models achieve milliseconds of inference speed. Additionally, the reasoning speed of the Port-LLM model is faster than that of the Prompt-Port-LLM model. This discrepancy can be attributed to the fact that the Prompt-Port-LLM model necessitates additional time to acquire dynamic prompt prior to each prediction, whereas the Port-LLM model, which utilizes LoRA fine-tuning, does not have this requirement.